OpenStudy (impracticaljoker):

Solve 7x − 2(x + 1) = 6x + 14. (5 points) −12 12 −16 16

OpenStudy (impracticaljoker):

@jhonyy9

OpenStudy (impracticaljoker):

@quickstudent

jhonyy9 (jhonyy9):

like a first step distribute these parentheses

OpenStudy (impracticaljoker):

Wouldn't that just be make 7x into 8x? That wasn't very descriptive.

jhonyy9 (jhonyy9):

no multiplie the 2 by x and by 1 inside parentheses

OpenStudy (quickstudent):

So here's the equation 7x - 2(x + 1) = 6x + 14 jhonyy9 can you let me explain this time? :)

jhonyy9 (jhonyy9):

ok. was my pleasure joker good luck bye bye

OpenStudy (impracticaljoker):

:)

OpenStudy (quickstudent):

7x - 2(x + 1) = 6x + 14 So yes, you do have to start by using the distributive property. Look: 7x - 2(x + 1) = 6x + 14 7x (-2 * x) + (-2 * 1) = 6x + 14 7x - 2x - 1 = 6x + 14 Get it?

OpenStudy (impracticaljoker):

Yes, distributing is kind of confusing but I get what you're saying

OpenStudy (impracticaljoker):

Next combine like terms right?

OpenStudy (quickstudent):

Yep:)

OpenStudy (impracticaljoker):

There are 3 x's this time. Do we make all those into one.

OpenStudy (quickstudent):

Woops, I made one mistake. Here's the correction: 7x - 2x - 2 = 6x + 14

OpenStudy (quickstudent):

You can only combine what is on one side of the equals sign. So the 7x and -2x can be combined, but not 6x.

OpenStudy (quickstudent):

@ImpracticalJoker ?

OpenStudy (impracticaljoker):

Sorry! My internet crashed

OpenStudy (quickstudent):

Oh, that's what I thought

OpenStudy (impracticaljoker):

How would we combine a positive and a negative?

OpenStudy (quickstudent):

It's simple. combining 7x and -2x is the same as doing 7x - 2x. what should it equal?

OpenStudy (impracticaljoker):

5

OpenStudy (impracticaljoker):

5x

OpenStudy (quickstudent):

Yes, so now we have 5x - 2 = 6x + 14

OpenStudy (impracticaljoker):

5x - 2 = 6x + 14

OpenStudy (quickstudent):

Yes! so do you know what to do now?

OpenStudy (quickstudent):

Looks like your internet is not working again:(. OK then, here's the rest for you to understand: 5x - 2 = 6x + 14 5x - 2 - 5x = 6x + 14 - 5x -2 = x + 14 -2 - 14 = x + 14 - 14 -16 = x Do you get it? If there's anything you see that you don't understand, just ask:)

OpenStudy (impracticaljoker):

I get it.My internet is acting up for some reason

OpenStudy (impracticaljoker):

Can you help with some more?

OpenStudy (quickstudent):

Sure, post it:)

OpenStudy (impracticaljoker):

How many solutions does the following equation have? |4x + 12| = 0 (5 points) No solution One solution Two solutions Infinitely many solutions

OpenStudy (quickstudent):

Hmm, sorry, I don't remember how to do this one:( Open this in a new question and I'll see if I can tag someone

OpenStudy (impracticaljoker):

ok